OpenStudy (anonymous):

A multiple-choice quiz has 15 questions, each with 4 possible answers of which only 1 is the correct answer. What is the probability that a student gets a. exactly 10 correct answers b. at least 1 correct answer c. 8 to 12 correct answer

OpenStudy (anonymous):

binomial distribution for each of these

OpenStudy (anonymous):

\[p=\frac{1}{4},1-p=\frac{3}{4}\] \[P(x=10)=\binom{15}{10}\left (\frac{1}{4}\right)^{10}\left(\frac{3}{4}\right)^5\]

OpenStudy (anonymous):

b) at least one means not none \[1-\left(\frac{3}{4}\right)^{15}\]

OpenStudy (anonymous):

8 to twelve has no shortcut, gotta add them all up

OpenStudy (anonymous):

add?

OpenStudy (anonymous):

c